OpenStudy (kaloemalia):

@butterflydreamer @babynini A line has a slope of 3 and contains the point (−1, −8). Which equations represent the line? Choose all answers that are correct. A. −3x + y = −5 B. 3x + y = −5 C. (y – 8) = 3(x − 1) D. (y + 8) = 3(x + 1) i think it is a and d?

Miracrown (miracrown):

Yes, A and D are correct

OpenStudy (babynini):

@Miracrown doesn't A make it a negative slope?

OpenStudy (babynini):

Oh sorry, nevermind!

Miracrown (miracrown):

Nope, as that is not in the slope-intercept form

OpenStudy (kaloemalia):

thanks guys!

OpenStudy (kaloemalia):

one more @Babynini @micracrown ?

OpenStudy (babynini):

Sure :) I shall try my very best!

OpenStudy (kaloemalia):

ok theres screen shots

OpenStudy (kaloemalia):

@Babynini

OpenStudy (kaloemalia):

theyre out of order

OpenStudy (kaloemalia):

@Babynini

OpenStudy (babynini):

Ok, give me a moment to open all those :P

OpenStudy (kaloemalia):

ok

OpenStudy (babynini):

The first thing to do would be write the equation in it's usual form.

OpenStudy (kaloemalia):

ok so.. that would be..?

OpenStudy (kaloemalia):

idk how to do that

OpenStudy (babynini):

the equation for a graph is: y =(slope)x + (y intercept) So simplify what you have: (y+1)=3/5(x-3) to fit that. Does that make sense?

OpenStudy (kaloemalia):

yes

OpenStudy (babynini):

Ok, work that out and then show it to me

OpenStudy (babynini):

(hint: it's helpful to make the "1" on the left have a denominator of 5 like the other fractions. So your new equation would be: y + 5/5=3/5(x-3)

OpenStudy (kaloemalia):

i dont really get it

OpenStudy (babynini):

Hm, okay. Let's see.

OpenStudy (babynini):